OpenStudy (anonymous):

WILL GIVE MEDAL IF YOU HELP Which choice describes a question that could be answered by moving the decimal point 3 places to the right? A. multiplying 31.04 by 100 B. multiplying 31.04 by 1,000 C. dividing 31.04 by 102 D. dividing 31.04 by 103

OpenStudy (anonymous):

I think it is A. but i am not sure

OpenStudy (freckles):

multiplying/dividing by 10^2 think 2 spots for example 4*10^2=400 or 4/10^2=.04 multiplying/dividing by 10^3 think 3 spots for example 4*10^3=4000 or 4/10^3=.004 So multiplying/dividing by 10^n think n spots

OpenStudy (freckles):

by the way you do know 10^3=1000 and 10^2=100 right?

OpenStudy (freckles):

and 10^4=10000 and 10^5=100000 and so on

OpenStudy (anonymous):

q